OpenStudy (anonymous):

a car traveling at +7.0 m/s accelerates at a rate of +0.80 m/s^2 for an interval of 2.0 s. findddd vf...

OpenStudy (anonymous):

vf=vi+at=7.0m/s+2s(0.80m/s^2)=8.6m/s

OpenStudy (anonymous):

\[v=u +at\] v= final velocity u= initial velocity a= acceleration t= time

OpenStudy (anonymous):

y=u+at, 7+(0.80*2) 7+1.6=8.6. u=initial velocity. v=final velocity. a=accelaration